Exam Item 3 Item 3 ho Part A Find the rms voltage across each element in an RLC circuit with R 99 kn.c 0.14 pF, and L 27 mH.The generator supplies an mms voltage of 110 V at a frequency of 65.0 Hz Express your answers using two significant figures. Enter your answers numerically separated by commas. vh vi, ve Submit My Answers Give Up Provide Feedback C nue MacBook
Answer #1

Z=R+jXL -jXC
R=9900 Ohm
XL=2*pi*f*L= 11.027 Ohm
XC = 1/(2*pi*f*C) = 17489.55 Ohm

Z=9900-j17478.52 Ohm
|Z| = 20087.53 Ohm

I= V/Z = 6.72mA

Vr = I*R = 66.5V
VL = I*XL = 0.074V
VC= I*XC = 117.53V